# Bitangent surfaces and involutions of quartic surfaces

### v2

@@ 5. Quartic surfaces: $p=2$
So, the image of $X$ is the intersection of two hypersurfaces of degree $4$ with the trivial canonical sheaf. It is singular along the $V(p_{0},p_{1},p_{4})$. The quotient $Y=X/(\sigma)$ is isomorphic to the normalization of this surface. Note that polynomials $p_{i}$ do not generate the algebra of invariants of the projective coordinate ring of $X$. In fact, $x_{0}(x_{3}^{3}+x_{0}x_{1})$ and $x_{1}(x_{0}^{3}+x_{1}^{2}x_{2})$ are invariant modulo $(F)$. The map $\phi:Y\to S(\sigma)$ is just the projection map given by $(p_{0},p_{1},p_{3})$. It blows down the lines $V(x_{0}+x_{2},x_{1}+x_{3})$ to the point $[0,0,1]$.
6. Kummer quartic surfaces in characteristic $2$: ordinary case
@@
Kummer quartic surfaces in characteristic 2 are divided into three classes according to curves of genus 2 being ordinary, 2-rank 1 or supersingular (e.g. Katsura–Kondō [13]). In this section, we discuss, the simplest case, an ordinary Kummer quartic surface. The Kummer quartic surface $X$ in characteristic $2$ associated with an ordinary genus 2 curve $C$ is given by
@@
where $a,b,c$ are non-zero constants. They coincide with the coefficients of the Igusa canonical model of $C$ (Laszlo–Pauly [16], [17], Katsura-Kondō [13]). The Kummer quartic surface $X$ has four singular points
@@
all of which are rational double points of type $D_{4}$ (Shioda [21], Katsura [12]), and $X$ has four tropes defined by the hyperplane sections $x=0,y=0,z=0,w=0$, respectively.
@@
It is enough to consider the first pair of skew lines. A line passing through a point $q=[x_{0},y_{0},z_{0},w_{0}]$ not on $\ell_{1}$ or $\ell_{2}$ is the intersection of two planes $\langle\ell_{1},q\rangle$ and $\langle\ell_{2},q\rangle$. Its parametric equation is
@@
It intersects $\ell_{1}$ at the point $[0,0,z_{0},w_{0}]$ and intersects $\ell_{2}$ at the point $[x_{0},y_{0},0,0]$. Plugging in these equations into (6.1), we get
@@
[…]
$[s:t]=[z_{0}w_{0}:x_{0}y_{0}],$
This defines an explicit bitangent involution $\sigma$ of $X$. We see that it is not defined only at singular points $p_{1},\ldots,p_{4}$ of $X$. It is obvious that the involution is the restriction of the Cremona involution
$T_{1}:[x,y,z,w]\dashrightarrow[xzw,yzw,xyz,xyw].$
It is equal to the product of the standard inversion transformation $T$ and the involution
@@
Note that $g_{1}$ is induced from a translation of $J(C)$ by a non-zero 2-torsion. There are three non-zero 2-torsions on $J(C)$ and all three bitangent involutions are the products of $T$ and the involutions induced from the translations.
@@
The surface $X$ admits three bitangent involution $\sigma_{1},\sigma_{2},\sigma_{3}$. The congruence of lines $S(\sigma_{i})$ is equal to the congruence of lines intersecting two skew lines from the previous lemma. The fixed curve $X^{\sigma_{i}}$ is an elliptic curve of degree $4$, and it is cut out set-theoretically by the quadric $V(xy+zw)$.
@@
It is enough to consider the first involution defined by the two lines $\ell_{1}=V(x,y)$ and $\ell_{2}=V(z,w)$. We denote it by $\sigma$. We have already proved the first assertion. Its fixed locus of $T_{1}$ is the quadric $Q=V(xy+zw)$. It cuts out $X$ along a complete intersection of two surfaces:
@@
Plugging in $y=zw/x$ in the second equation, we find that the intersection is a quartic curve taken with multiplicity $2$. It passes through the singular points, and these points are nonsingular. We also check (substituting $x=1$) that its projection to the plane is a smooth cubic curve. Thus, $C=X^{\sigma}$ is, set-theoretically, a smooth quartic elliptic curve.
@@
We consider only the pencil defined by $\ell_{1}^{\perp}=V(y+tx)$ with parameter $t\in\mathbb{P}^{1}$. We check that, for a general parameter $t$, the plane quartic $S(x,tx,z,w)$ has two cusps at $p_{3}$ and $p_{4}$, hence its geometric genus is equal to $1$. The cusps are the base points of the pencil of quartics. There are special values of the parameter $t$: $0,\infty,a/b,c/b$. For $t=0$ (resp. $t=\infty$), the quartic is a double conic, it passes through the third singular point $p_{1}\in\ell_{2}$ (resp. $p_{2}\in\ell_{2}$). It is one of the four trope-conics on $X$, i.e. conics cut out set-theoretically by one of trope-planes:
$C_{123}$ $=$ $V(axy+bxz+cyz),$
$C_{124}$ $=$ $V(axy+byw+cxw),$
$C_{134}$ $=$ $V(azw+bxz+cxw),$
$C_{234}$ $=$ $V(azw+byw+cyz).$
Let $E_{i}=E_{0}^{(i)}+E_{1}^{(i)}+E_{2}^{(i)}+E_{3}^{(i)},i=1,2,3,4,$ be the exceptional curves over singular points $p_{i}$ of $X$, where $E_{0}^{(i)}$ is the central component. We check that the proper transform of $V(y)$ intersects one of the components $E_{i}^{(j)},i\neq 0,j=1,3,4$. The corresponding fiber of $|F_{1}|$ is of type $\tilde{D}_{6}$. If $t=a/b$, the quartic has one cusp at one of the base points and a cusp followed by infinitely near node at another base points. It defines a fiber of type $\tilde{A}_{1}^{*}$. ∎
@@
The involution $\tilde{\sigma}$ interchanges the proper transforms $\tilde{C}_{123}$ of $C_{123}$ with the central component $E_{0}^{(3)}$, and similarly for other trope conics. It acts on the corresponding reducible fiber via the symmetry of its dual graph.
@@
Let $|\mathfrak{f}_{1}|$ and $|\mathfrak{f}_{2}|$ be the rulings of the quadric $Q=S(\sigma)$ corresponding to the family of planes $\ell_{1}^{\perp}$ and $\ell_{2}^{\perp}$. Their pre-images under $\tilde{X}\to\tilde{Y}\to Q$ are the two elliptic pencils from Proposition 6.2. Let $L_{0}^{(1)},L_{\infty}^{(1)}$ (resp. $L_{0}^{(2)},L_{\infty}^{(2)}$) be the lines from $|\mathfrak{f}_{1}|$ (resp. $|\mathfrak{f}_{2}|$) corresponding to the planes $V(x),V(y)\in\ell_{1}^{\perp}$ (resp. $V(z),V(w)\in\ell_{2}^{\perp}$). Their pre-images in $\tilde{X}$ are the reducible fibers of type $\tilde{D}_{6}$ of $|F_{1}|$ (resp. $|F_{2}|)$. The image of the curve $\tilde{X}^{\tilde{\sigma}}$ in $Q$ is a smooth curve $B$ of bidegree $(2,2)$ that passes through the vertices of the quadrangle of lines $L=L_{0}^{(1)}+L_{\infty}^{(1)}+L_{0}^{(2)}+L_{\infty}^{(2)}$.
We leave it to the reader to check the following assertion:
@@
The morphism $\tilde{\phi}_{\sigma}:\tilde{X}/(\tilde{\sigma})\to S(\sigma)$ is the blow-up of the eight points on $S(\sigma)$: four vertices of the quadrangle of lines $L$ and infinitely near points to them corresponding to the tangent direction of the curve $B$. The pre-images of the two generators in $|\mathfrak{f}_{i}|$ corresponding to the ramification points of the projection map $B\to|\mathfrak{f}_{i}|^{*}\cong\mathbb{P}^{1}$ are reducible fibers of $|F_{i}|$ of type $\tilde{A}_{1}^{*}$.
So far, we have found that the bitangent surface ${\mathrm{Bit}}(X)$ contains three irreducible components of bidegree $(m,n)=(1,1)$. They correspond to three bitangent involutions $\sigma_{i}$. It also contains four irreducible components of bidegree $(0,1)$ corresponding to the trope-conics $C_{ijk}$.
@@
The surface ${\mathrm{Bit}}(X)$ is of bidegree $(m,n)=(3,7)$ in $G_{1}(\mathbb{P}^{3})$. It consists of $7$ irreducible components listed in above.
@@
It is enough to show that the class $n$ of ${\mathrm{Bit}}(X)$ is equal to $3+4=7$. This is the number of bitangent lines contained in a general plane in $\mathbb{P}^{3}$. We have already shown that $n\geq 7$. On the other hand, it is known that the number of bitangent lines of a smooth plane quartic in characteristic $2$ is $7,4,2$ or $1$ if the Hasse-Witt invariant is equal to $3,2,1$ or 00, respectively (Resmini [19], Stöhr and Voloch [22]). The assertion now follows. ∎
@@
of $\mathbb{P}^{3}$. It restricts to a fixed-point-free involution $\sigma$ of $\tilde{X}$ with $\tilde{X}/(\sigma)$ isomorphic to an Enriques surface. Let us consider the Plücker embedding of the corresponding congruence of lines. The six minors $p_{ij}$ of the matrix
@@
The line $\ell$ is given by
$x_{0}(y_{0}^{2}+z_{0}^{2})x+y_{0}(x_{0}^{2}+z_{0}^{2})y+z_{0}(x_{0}^{2}+y_{0}^{2})z=x_{0}(y_{0}^{2}+w_{0}^{2})x+y_{0}(x_{0}^{2}+w_{0}^{2})y+w_{0}(x_{0}^{2}+y_{0}^{2})w=0.$
By putting these equations into (6.1), we get
$xy((x_{0}^{2}(y_{0}^{2}+z_{0}^{2})(y_{0}^{2}+w_{0}^{2})/z_{0}w_{0}(x_{0}^{2}+y_{0}^{2})^{2})x^{2}+(y_{0}^{2}(x_{0}^{2}+z_{0}^{2})(x_{0}^{2}+w_{0}^{2})/z_{0}w_{0}(x_{0}^{2}+y_{0}^{2})^{2})y^{2})$
$\equiv xyzw\ ({\rm modulo\ squares}).$
In general, the coefficients of $x^{3}y,xy^{3}$ are not equal to zero, and hence, the assertion holds. ∎
Let $\ell$ be a line passing through $p_{0}$ and $\sigma(p_{0})$. Let $\ell\cap X=\{p_{0},\sigma(p_{0}),p_{1},p_{2}\}$. The image of $\ell$ under $\sigma$ is a cubic curve, and hence, $p_{1},p_{2}$ are not conjugate to $p_{0},\sigma(p_{0})$. Thus, the map $X\dasharrow G_{1}(\mathbb{P}^{3})$ sending $p_{0}$ to $\ell=\langle p_{0},\sigma(p_{0})\rangle$ has degree 2 onto its image. Therefore, the Enriques surface $\tilde{X}/(\sigma)$ can be embedded into $G_{1}(\mathbb{P}^{3})$ satisfying a cubic relation (6.3). This suggests the following question:
@@
7. Kummer quartic surfaces in $p=2$: $2$-rank $1$ case

### v3

@@ 5. Quartic surfaces: $p=2$
Therefore, the image of $X$ in $\mathbb{P}^{3}/(\sigma)$ is the intersection of two hypersurfaces of degree $4$, so it has trivial dualizing sheaf. It is singular along the line $V(p_{0},p_{1},p_{4})$. The quotient $X/(\sigma)$ is isomorphic to the normalization of this surface. Note that the polynomials $p_{i}$ do not generate the algebra of invariants of the projective coordinate ring of $X$. In fact, $x_{0}(x_{3}^{3}+x_{0}x_{1})$ and $x_{1}(x_{0}^{3}+x_{1}^{2}x_{2})$ are invariant modulo $(F)$.
The map $\phi:X/(\sigma)\to S(\sigma)$ is just the projection map given by $(p_{0},p_{1},p_{3})$. It shows that the congruence of lines $S(\sigma)$ is isomorphic to $\mathbb{P}(1,1,2)$. The image $\phi(\ell)$ of the line $\ell$ is the singular point $[0,0,1]$ of $\mathbb{P}(1,1,2)$. The images of the four invariant lines are two nonsingular points.
Lemma 5.7.
With no assumption on the characteristic, suppose ${\mathrm{Bit}}(X)$ contains a $\beta$-plane of lines in a plane $\Pi=V(L)$, where $L$ is a linear form. Then, the equation of $X$ can be written in the form
(5.1) $Q(x,y,z,w)^{2}+L(x,y,z,w)F(x,y,z,w)=0,$
where $Q$ is a quadratic form and $F$ is a cubic form. The singular locus of $X$ is equal to $V(L,Q,F)\cup V(Q,\textrm{Sing}(V(F))$. In particular, $X$ is a singular quartic surface.
Proof.
Let $C=X\cap\Pi$. Any line in $\Pi$ is a bitangent of $X$. Thus, it is bitangent to $C$. This could happen only if $C$ is a double conic, reducible or not. Thus, $\Pi$ is a trope-conic, and its equation can be written as in (5.1).
Without loss of generality, we may assume that $L=x$ is a coordinate plane. Taking the partials, we find the singular locus is indeed $V(x,Q,F)$. ∎
Recall that the $p$-rank of a smooth curve $C$ of genus $g>0$ over an algebraically closed field of characteristic $p>0$ is the $p$-rank of the elementary abelian group $J(C)[p]$ of $p$-torsion points of its jacobian variety of $C$. It takes values in the set $[0,g]$. When $p=2$ and $C$ is a smooth plane quartic, the $p$-rank takes values in $\{3,2,1,0\}$ and can be characterized by the number of bitangents of $C$ equal to $7,4,2,1$, respectively [18], [22, §3].
The proof of the following proposition can be found in [26, Proposition 1] (see also [18]).
Proposition 5.8.
Let $C$ be a smooth plane quartic curve over an algebraically closed field of characteristic $2$. Then, it is projectively equivalent to one of the curves
  1. (1)
    $Q^{2}+xyz(x+y+z)=0,\quad Q(1,0,0),\ Q(0,1,0),\ Q(0,0,1),\ Q(1,1,0),\ Q(1,0,1),\\ Q(0,1,1),\ Q(1,1,1)\neq 0$;
  2. (2)
    $Q(x,y,z)^{2}+xyz(y+z)=0,\quad Q(1,0,0),\ Q(0,1,0),\ Q(0,0,1),\ Q(0,1,1)\neq 0$;
  3. (3)
    $Q(x,y,z)^{2}+xy(y^{2}+xz)=0,\quad Q(1,0,0),\ Q(0,0,1)\neq 0$;
  4. (4)
    $Q(x,y,z)^{2}+x(y^{3}+x^{2}z)=0,\quad Q(0,0,1)\neq 0$.
One check that the number of bitangents is indeed equal to $7,4,2,1$, respectively. More precisely, the bitangents are
  1. (1)
    $V(x),V(y),V(z),V(x+y),V(y+z),V(z+x),V(x+y+z).$
  2. (2)
    $V(x),V(y),V(z),V(y+z)$;
  3. (3)
    $V(x),V(y)$;
  4. (4)
    $V(x)$.
Computing the stabilizer group of projective transformations that leave the normal form (1)–(4) invariant, we see that smooth quartics of $2$-rank less than $3$ form three hypersurfaces in the space of all smooth plane quartics.
It is clear that there exist smooth quartic surfaces with a general plane section isomorphic to a plane quartic curve with $p$-rank equal to $3$, we do not know whether it is true for plane curves of rank 2 and 1. The next proposition shows that it is not true for supersingular plane quartic curves.
Proposition 5.9.
Let $X$ be a normal quartic surface whose general hyperplane section is isomorphic to a plane quartic with the $2$-rank equal to 00. Then, ${\mathrm{Bit}}(X)$ contains a $\beta$-plane, and hence, it is singular.
Proof.
Let
$F:=\{(\ell,\Pi)\in{\mathrm{Bit}}(X)\times\check{\mathbb{P}}^{3}:\ell\subset\Pi\},$
and let $f:F\to\check{\mathbb{P}}^{3}$ be the second projection. Consider a general plane section as a plane curve over a field of rational functions of the dual projective space ˇ$\check{\mathbb{P}}^{3}$. Applying Proposition 5.8, we find a family $f^{\prime}:F^{\prime}\to B$ where $B$ is a finite cover of ˇ$\check{\mathbb{P}}^{3}$ such that $F^{\prime}$ contains an irreducible component that is mapped to a point. This corresponds to the plane $V(x)$ in equation (4) from Proposition 5.8. The image of this component in $F$ is a $\beta$-plane $\Pi$ contained in ${\mathrm{Bit}}(X)$. Further computation shows that there is nothing else. ∎
Example 5.10.
(suggested by the referee). Consider the surface $X$ with the equation
$w^{2}(ax+by+z)^{2}+x(y^{3}+x^{2}z)=0.$
The surface $X$ has a unique (non-rational) singular point $[0,0,0,1]$ and an ordinary node $[0,0,1,0]$. The point $[0,0,0,1]$ is a unique inseparable projection center of $X$. In fact, take a general point $P=[x_{0},y_{0},z_{0},w_{0}]$ in $\mathbb{P}^{3}$. Substituting the parametric equation $[x,y,z,w]=[sx_{0}+tu_{0},sy_{0}+tu_{1},sz_{0}+tu_{2},sw_{0}+tu_{3}]$, we find one bitangent line that connects $P$ with the point $[0,0,0,1]$. Thus, ${\mathrm{Bit}}(X)$ contains an $\alpha$-plane. On the other hand, a general hyperplane section $w=\alpha x+\beta y+\gamma z$ is the case of Proposition 5.8, that is, it is a plane quartic curve with the 2-rank 0. By Proposition 5.9, it contains a $\beta$-plane. Thus, ${\mathrm{Bit}}(X)$ contains the union of an $\alpha$-plane and a $\beta$-plane. Further computation shows that there is nothing else.
6. Kummer quartic surfaces in characteristic $2$, ordinary case
@@
Kummer quartic surfaces in characteristic 2 are divided into three classes according to curves of genus 2 being ordinary, 2-rank 1, or supersingular (see, e.g., [13]). In this section, we discuss the simplest case, an ordinary Kummer quartic surface. The Kummer quartic surface $X$ in characteristic $2$ associated with an ordinary genus 2 curve $C$ is given by
@@
where $a,b,c$ are non-zero constants. They coincide with the coefficients of the Igusa canonical model of $C$ [16], [17], [13]. The Kummer quartic surface $X$ has four singular points
@@
all of which are rational double points of type $D_{4}$ [21], [12], and $X$ has four tropes defined by the hyperplane sections $x=0,y=0,z=0,w=0$, respectively.
Applying Proposition 5.8, we find that a general hyperplane section of $X$ is a plane quartic with the $2$-rank equal to $3$. So, we expect that the class of $X$ is equal to $7$. Computing the partial derivatives of the polynomial defining $X$, we find that there are no inseparable projection centers of $X$. So, ${\mathrm{Bit}}(X)$ does not contain $\alpha$-planes.
@@
It is enough to consider the first pair of skew lines. A line $\ell$ passing through a point $q=[x_{0},y_{0},z_{0},w_{0}]$ not on $\ell_{1}$ or $\ell_{2}$ is the intersection of two planes $\langle\ell_{1},q\rangle$ and $\langle\ell_{2},q\rangle$. The parametric equation of $\ell$ is
@@
The line $\ell$ intersects $\ell_{1}$ at the point $[0,0,z_{0},w_{0}]$ and intersects $\ell_{2}$ at the point $[x_{0},y_{0},0,0]$. Plugging in these equations into (6.1), we get
@@
We keep the notation in the proof of Lemma 6.1. If $q$ belongs to $X$, then it corresponds to $[s,t]=[1,1]$, so $[s,t]=[1,1]$ is one of the solutions of (6.2). The second solution is $[s,t]=[z_{0}w_{0},x_{0}y_{0}]$. This defines an explicit bitangent involution $\sigma_{1}$ of $X$. We see that it is not defined only at the singular points $p_{1},\ldots,p_{4}$ of $X$. The involution is the restriction of the Cremona involution
$T_{1}:[x,y,z,w]\dashrightarrow[xzw,yzw,xyz,xyw]$
which is equal to the composition of the standard inversion transformation $T$ and the involution
@@
Note that $g_{1}$ is induced from a translation of $J(C)$ by a non-zero 2-torsion point. There are three non-zero 2-torsion points on $J(C)$ and all three bitangent involutions are the products of $T$ and the involutions induced from the translations.
@@
The surface $X$ admits three bitangent involutions $\sigma_{1},\sigma_{2},\sigma_{3}$. The congruence of lines $S(\sigma_{i})$ is equal to the congruence of lines intersecting two skew lines from the previous lemma. The fixed curve $X^{\sigma_{i}}$ is an elliptic curve of degree $4$, and it is cut out set-theoretically by the quadric $V(xy+zw)$.
@@
It is enough to consider the first involution defined by the two lines $\ell_{1}=V(x,y)$ and $\ell_{2}=V(z,w)$. We denote it by $\sigma$. We have already proved the first assertion. The fixed locus of $T_{1}$ is the quadric $Q=V(xy+zw)$. The quadric intersects $X$ along a complete intersection of two surfaces:
@@
Plugging in $y=zw/x$ in the second equation, we find that the intersection is a quartic curve taken with multiplicity $2$. It passes through the singular points, and these points are nonsingular on the quartic. We also check (substituting $x=1$) that its projection to the plane is a smooth cubic curve. Thus, $C=X^{\sigma}$ is, set-theoretically, a smooth quartic elliptic curve.
@@
We consider only the pencil defined by $\ell_{1}^{\perp}=V(y+tx)$ with parameter $t\in\mathbb{P}^{1}$. We check that, for a general parameter $t$, the plane quartic $X\cap V(y+tx)$ has two cusps at $p_{3}$ and $p_{4}$, hence its geometric genus is equal to $1$. The cusps are the base points of the pencil of quartics. There are special values of the parameter $t$: $0,\infty,\sqrt{c/b},\sqrt{b/c}$. For $t=0$ (resp. $t=\infty$), the quartic is a double conic, it passes through the third singular point $p_{1}\in\ell_{2}$ (resp. $p_{2}\in\ell_{2}$). The conic is one of the four trope-conics on $X$, i.e. conics cut out set-theoretically by one of the trope planes:
$C_{123}$ $=$ $V(\sqrt{a}xy+\sqrt{b}xz+\sqrt{c}yz),$
$C_{124}$ $=$ $V(\sqrt{a}xy+\sqrt{b}yw+\sqrt{c}xw),$
$C_{134}$ $=$ $V(\sqrt{a}zw+\sqrt{b}xz+\sqrt{c}xw),$
$C_{234}$ $=$ $V(\sqrt{a}zw+\sqrt{b}yw+\sqrt{c}yz).$
Let $E_{i}=E_{0}^{(i)}+E_{1}^{(i)}+E_{2}^{(i)}+E_{3}^{(i)},i=1,2,3,4,$ be the exceptional curves over the singular points $p_{i}$ of $X$, where $E_{0}^{(i)}$ is the central component. We check that the proper transform of $V(y)$ intersects one of the components $E_{i}^{(j)},i\neq 0,j=1,3,4$. The corresponding fiber of $|F_{1}|$ is of type $\tilde{D}_{6}$. If $t=a/b$, the quartic has one cusp at one of the base points and a cusp followed by infinitely near node at another base points. The corresponding fiber of type $\tilde{A}_{1}^{*}$. ∎
@@
The involution $\tilde{\sigma}$ interchanges the proper transform $\tilde{C}_{123}$ of $C_{123}$ with the central component $E_{0}^{(3)}$, and similarly for other trope conics. It acts on the corresponding reducible fiber via a symmetry of its dual graph.
@@
Let $|\mathfrak{f}_{1}|$ and $|\mathfrak{f}_{2}|$ be the rulings of the quadric $Q=S(\sigma)$ corresponding to the family of planes $\ell_{1}^{\perp}$ and $\ell_{2}^{\perp}$. Their pre-images under $\tilde{X}\to\tilde{Y}\to Q$ are the two elliptic pencils from Lemma 6.3. Let $L_{0}^{(1)},L_{\infty}^{(1)}$ (resp. $L_{0}^{(2)},L_{\infty}^{(2)}$) be the lines from $|\mathfrak{f}_{1}|$ (resp. $|\mathfrak{f}_{2}|$) corresponding to the planes $V(x),V(y)\in\ell_{1}^{\perp}$ (resp. $V(z),V(w)\in\ell_{2}^{\perp}$). Their pre-images in $\tilde{X}$ are the reducible fibers of type $\tilde{D}_{6}$ of $|F_{1}|$ (resp. $|F_{2}|)$. The image of the curve $\tilde{X}^{\tilde{\sigma}}$ in $Q$ is a smooth curve $B$ of bidegree $(2,2)$ that passes through the vertices of the quadrangle of lines $L=L_{0}^{(1)}+L_{\infty}^{(1)}+L_{0}^{(2)}+L_{\infty}^{(2)}$.
We summarize our discussion above with the following assertion:
@@
The morphism $\tilde{\phi}_{\sigma}:\tilde{X}/(\tilde{\sigma})\to S(\sigma)$ is the blow-up of the following eight points on $S(\sigma)$: four vertices of the quadrangle of lines $L$ and infinitely near points to them corresponding to the tangent direction of the curve $B$. The pre-images of the two fibers in $|\mathfrak{f}_{i}|$ corresponding to the ramification points of the projection map $B\to|\mathfrak{f}_{i}|^{*}\cong\mathbb{P}^{1}$ are reducible fibers of $|F_{i}|$ of type $\tilde{A}_{1}^{*}$.
So far, we have found that the bitangent surface ${\mathrm{Bit}}(X)$ contains three irreducible components of bidegree $(m,n)=(1,1)$. They correspond to three bitangent involutions $\sigma_{i}$. The surface also contains four irreducible components of bidegree $(0,1)$ corresponding to the trope-conics $C_{ijk}$.
@@
Let $X$ be the Kummer quartic surface $X$ in characteristic $2$ associated with an ordinary curve of genus $2$. Then, the surface ${\mathrm{Bit}}(X)$ is of bidegree $(m,n)=(3,7)$ in $G_{1}(\mathbb{P}^{3})$ which consists of $7$ irreducible components, three of bidegree $(1,1)$ corresponding to three bitangent involutions and four of bidegree $(0,1)$ corresponding to four tropes.
@@
For any plane $\Pi$ containing $p$, the plane quartic curve $X\cap\Pi$ has bitangent lines forming a line in $G_{1}(\mathbb{P}^{3})$, which implies that the plane quartic curve is a double conic. This is impossible for a normal quartic $X$. As we mentioned before, Lemma 6.1, ${\mathrm{Bit}}(X)$ does not contain $\alpha$-planes. Thus it is enough to show that the class $n$ of ${\mathrm{Bit}}(X)$ is equal to $3+4=7$. This is the number of bitangent lines contained in a general plane in $\mathbb{P}^{3}$. We have already shown that $n\geq 7$. On the other hand, it is known that the number of bitangent lines of a smooth plane quartic in characteristic $2$ is $7,4,2$ or $1$ if the Hasse-Witt invariant is equal to $3,2,1$ or 00, respectively [22]. The assertion now follows. ∎
@@
of $\mathbb{P}^{3}$. The restriction of $T$ to $\tilde{X}$ is a fixed-point-free involution $\sigma$ with $\tilde{X}/(\sigma)$ isomorphic to an Enriques surface. Let us consider the Plücker embedding of the corresponding congruence of lines. The six minors $p_{ij}$ of the matrix
@@
The following proof is suggested by a referee. By plugging the parametric equation
$[x_{0}+ty_{0}z_{0}w_{0},y_{0}+tx_{0}z_{0}w_{0},z_{0}+tx_{0}y_{0}w_{0},w_{0}+tx_{0}y_{0}z_{0}]$
of the line $\ell$ into the equation (6.1) of $X$, we can see that the term which is linear in $t$ is equal to $(x_{0}+y_{0}+z_{0}+w_{0})t$. Since a general point does not lie in the hyperplane $V(x+y+z+w)$, $\ell$ is not a bitangent line. ∎
Let $\ell$ be a line passing through $p_{0}$ and $\sigma(p_{0})$. Let $\ell\cap X=\{p_{0},\sigma(p_{0}),p_{0}^{\prime},p_{0}^{\prime\prime}\}$. The image of $\ell$ under $\sigma$ is a cubic curve, and hence, $p_{0}^{\prime},p_{0}^{\prime\prime}$ are not conjugate to $p_{0},\sigma(p_{0})$ under the action of $\sigma$. Thus, the map $X\dasharrow G_{1}(\mathbb{P}^{3})$ sending $p_{0}$ to $\ell=\langle p_{0},\sigma(p_{0})\rangle$ has degree 2 onto its image. Therefore, the Enriques surface $\tilde{X}/(\sigma)$ can be embedded into $G_{1}(\mathbb{P}^{3})$ satisfying a cubic relation (6.3). This suggests the following question:
@@
7. Kummer quartic surfaces in $p=2$, $2$-rank $1$ case